# SynthAorta: A 3D Mesh Dataset of Parametrized Physiological Healthy Aortas

### v1

@@ Abstract
Keywords: Aorta modeling, aorta parametrization, convolution surfaces, mesh generation, synthetic models
@@ 1 Introduction
Computational methods for the modeling of the vascular system have been growing in popularity in recent years, as they offer a plethora of possibilities for clinicians and engineers alike. A non-exhaustive list includes modeling of digital twins or virtual cohorts, supporting physicians in making assessments and predictions, as well as analyzing the effects of various parameters on target mechanical quantities, and consequentially on the health of the patients. An aspect of particular interest, which indisputably has a crucial impact on the mechanics of the aorta, is its geometry. Many works have considered this topic [34, 28, 15, 46, 24, 37]. However, a barrier that is ever-present lies in obtaining usable patient-specific models. Existing works usually make use of a small number of patient-specific geometries. This is completely understandable when one considers the difficulties and the user interaction required for obtaining computational models from patient-specific CT scans [51, 23, 13, 20]. Besides, obtaining the CT scans themselves already presents a challenge [2, 52].
@@
Imposing a parametrized model to generate (smooth) aortic geometries presents a challenging task. In our approach, we only require a single base model, containing a centerline which consists only of points and straight segments. Then, the aforementioned parameters are applied only to the centerline, which is an extremely simple object. While it is clear that a centerline may be extracted from the surface of a patient-specific aorta, going the other way around may not seem like a viable task. Here, we make use of convolution surfaces [10, 21], a graphics methodology to render surfaces based purely on centerline and radial data. Essentially, a small number of points, segments, and associated radii (order of magnitude of $20-30$) are enough to describe a smooth surface of an aorta. Thus, parametrizing the centerline and radius information enables the parametrization of the entire aorta.
Even though the generated surfaces are useful, they are still insufficient to perform numerical simulations. The primary goal herein is to enable users to apply methods based on the standard three-dimensional finite-element method (FEM). With this objective in mind, it is necessary to produce volumetric discretizations of the generated aortic surfaces, i.e., meshes. The standard methodology for non-trivial domains such as aortas is unstructured meshing [31]. This is justified by the flexibility and robustness of the meshing procedure, as well as a multitude of off-the-shelf software tools [22, 45]. On the other hand, it is often worth investing extra effort into the generation of structured meshes [3, 13], as they offer certain advantages. Examples include exact control of the element count, straightforward formation of boundary layers, flow-oriented elements, and finally, the control of the local mesh structure. Hence, the resulting meshes have the same node placement independently of the geometry variation, greatly boosting their applicability for various applications such as machine learning or reduced order modeling. To realize the mesh generation, we make use of the block-structured approach presented in [13], extended in [12, 14]. It fits well with the approach described herein since the required input is exactly a convolution surface description of the domain.
@@ 2.1 Convolution surfaces
where $\Gamma:[0,1]\rightarrow\mathbb{R}^{3}$ denotes the parametrization of a centerline segment $[\mathbf{a},\,\mathbf{b}]$ defined as
$\Gamma_{S}(s)=(1-s)\mathbf{a}+s\mathbf{b}.$ (2)
The compactly supported kernel function $K:\mathbb{R}\rightarrow\mathbb{R}$ is defined as
@@
and the distance function $g:\mathbb{R}^{3}\times\mathbb{R}^{3}\rightarrow\mathbb{R}$ as
$g(\mathbf{x},\mathbf{y})=\sqrt{\mathbf{x}^{T}\,G(\mathbf{y})\,\mathbf{x}}.$ (4)
The matrix $G$ is computed from the frame matrix $U$ (constant along a segment, as we do not use curved centerlines) and the diagonal matrix $D$, which embeds the radius information:
$G=UDU^{T}.$ (5)
@@ 2.2 Statistical geometric parametrization
Table 1: Aortic parameters and their probabilistic distributions. The parameters in group A are purposefully modified to create new aortic datasets. The parameters in group B are produced as a result of the newly generated aorta. The distribution parameters, e.g., $(a,b)$, are defined as mean and variance for a normal distribution, as location and scale for a Gumbel distribution, and as minimum and maximum values for a uniform distribution.
Group Parameter Notation Collected distribution Unit Reference
type parameters
A Ascending aorta radius $r_{\text{AA}}$ Normal $(14.08,2.59)$ mm [36, 5, 16, 50]
Aortic arch radius $r_{\text{Arch}}$ Normal $(13.85,3.16)$ mm [36, 16]
Descending aorta radius $r_{\text{DA}}$ Normal $(12.17,2.31)$ mm [49, 36, 16]
Curvature radius multiplier $\gamma_{\text{c}}$ Normal $(1.07,0.01)$- [28, 11, 37]
Descending aorta adjustment $\rho$-
Nonplanar Fourier multiplier $\delta_{3}$-
Nonplanar Fourier multiplier $\delta_{4}$-
B Brachiocephalic artery (BCA) $r_{\mathrm{BCA}}$ Normal $(6.10,0.12)$ mm [33]
Common carotid artery (CCA) $r_{\mathrm{CCA}}$ Normal $(3.26,0.03)$ mm [43]
Left subclavian artery (LSA) $r_{\mathrm{LSA}}$ Normal $(5.42,0.07)$ mm [6]
Centerline curvature radius $R_{\text{c}}$ Normal $(39.30,3.92)$ mm [28, 11, 37]
Table 2: Parameter distributions in the actual dataset. Some of the parameter combinations produce invalid or unphysiological domains, causing the convolution surface to blend, in particular between the carotid arteries. In a minority of the cases, the meshing algorithm produces invalid elements. Thus, the final distributions are slightly affected.
Group Parameter Notation SynthAorta distribution Unit
type parameters
A Ascending aorta radius $r_{\text{AA}}$ Normal $(13.69,2.30)$ mm
Aortic arch radius $r_{\text{Arch}}$ Normal $(13.00,1.99)$ mm
Descending aorta radius $r_{\text{DA}}$ Normal $(12.17,2.31)$ mm
Curvature radius multiplier $\gamma_{\text{c}}$ Gumbel $(1.10,0.06)$-
Descending aorta adjustment $\rho$ Uniform $(0,1)$-
Nonplanar Fourier multiplier $\delta_{3}$ Normal $(1,0.09)$-
Nonplanar Fourier multiplier $\delta_{4}$ Normal $(1,0.09)$-
B Brachiocephalic artery (BCA) $r_{\mathrm{BCA}}$ Normal $(6.10,0.12)$ mm
Common carotid artery (CCA) $r_{\mathrm{CCA}}$ Normal $(3.26,0.03)$ mm
Left subclavian artery (LSA) $r_{\mathrm{LSA}}$ Normal $(5.42,0.07)$ mm
Centerline curvature radius $R_{\text{c}}$ Gumbel $(40.41,2.40)$ mm
@@ 2.2.1 Aortic radii
We use a Gaussian random field, using a Karhunen-Loève (KL) expansion, to create realistic radii samples of the three aortic segments. This method helps us generate smooth, interconnected variations along the length of the aorta [18]. The correlation structure is manually adjusted to reflect the physiological connections between the different segments of the aorta, ensuring that the produced aortas maintain realistic anatomical characteristics. We determine the average and variability of the random field from weighted values obtained from medical literature. These variations introduce physiological differences within each realization. The mean and variance of the random field are defined as $\mu_{\text{w}}$ and $\sigma_{\text{w}}$ and are derived from weighted values reported in the medical literature, see Table 1.
@@ 2.2.3 Centerline curvature radius
For the rest of the section, we assume the base centerline has been mapped to a single best-fitting plane using a typical SVD-based procedure. We define the point $\mathbf{A}$ in the ascending aorta, and $\mathbf{B}$ as the point on the centerline which is on the same axial plane as $\mathbf{A}$ following [28] and [37], shown in Fig. 3a.
@@
Denoting the centerline points of the main aorta which lie between them: $\mathbf{A}=\mathbf{x_{1}},\mathbf{x_{2}},\dots,\mathbf{x_{n-1}},\mathbf{x_{n}}=\mathbf{B}$, the centerline curvature radius is determined as follows. For the initial patient-specific centerline, we determine the center point $\mathbf{x_{\text{c}}}=(x_{\text{c}},y_{\text{c}})$ and the radius $r$ of the circle by solving the following optimization problem
$\underset{\mathbf{x_{\text{c}}},r}{\min}\,\,{\sum_{i=1}^{n}\left(\left\|\mathbf{x_{i}}-\mathbf{x_{\text{c}}}\right\|^{2}-r^{2}\right)^{2}},$ (8)
@@ 2.2.4 Descending aorta adjustment
Let us consider the centerline points $\mathbf{x_{n}},\,\dots,\mathbf{x_{n+k}}$, starting from $\mathbf{x_{n}}=\mathbf{B}$ and ending at the main outlet point of the centerline $\mathbf{x_{n+k}}=\mathbf{x_{\text{outlet}}}$, with the connections between neighbouring points $\mathbf{x_{j}}$ and $\mathbf{x_{j+1}}$. Let $\Delta\mathbf{B}\in\mathbb{R}$ be the displacement of the point $\mathbf{B}$ resulting from a change in the centerline curvature radius $R_{\text{c}}$. We introduce a new parameter $\rho$ uniformly distributed between 00 and $1$, defining the displacement at the outlet point $\mathbf{x_{\text{outlet}}}$ as
@@
$\Delta\mathbf{x_{j}}=\rho\frac{\|\mathbf{B}-\mathbf{x_{j}}\|}{\|\mathbf{B}-\mathbf{x_{\text{outlet}}}\|}\Delta\mathbf{B},\quad j=n+1,\dots,n+k-1.$ (14)
The displacement at $\mathbf{x_{j}}$ is a therefore a distance-based interpolation of displacements at $\mathbf{B}$ and $\mathbf{x_{\text{outlet}}}$. An example is illustrated in Fig. 3c.
@@ 2.2.5 Nonplanar displacement
$\mathbf{d}_{N}(\mathbf{x})=A_{0}+\sum_{n=1}^{N}\left(A_{n}\cos\left(nw\mathbf{x}\right)+B_{n}\sin\left(nw\mathbf{x}\right)\right),$ (15)
@@
$\hat{\mathbf{d}}_{2}^{(k)}(\mathbf{x})=A_{0}$ $+A_{1}\gamma_{1}\cos(w\mathbf{x})+B_{1}\gamma_{2}\sin(w\mathbf{x})$
$+A_{2}\delta_{3}\cos(2w\mathbf{x})+B_{2}\delta_{4}\sin(2w\mathbf{x}).$ (16)
@@ 2.3 Mesh generation
The algorithm builds a block-structure [1], a coarse domain subdivision, based on the input centerline. The block-structure generation is based on various centerline configurations occurring in the aorta. Then, the implicit convolution surface and the block-structure are combined to generate the mesh by using transfinite maps [53].

### v2

@@ Abstract
{IEEEkeywords}
Aorta modeling, aorta parametrization, convolution surfaces, mesh generation, synthetic models
@@ 1 Introduction
[…]
Computational methods for the modeling of the vascular system have been growing in popularity in recent years, as they offer a plethora of possibilities for clinicians and engineers alike. A non-exhaustive list includes modeling of digital twins or virtual cohorts, supporting physicians in making assessments and predictions, as well as analyzing the effects of various parameters on target mechanical quantities, and consequentially on the health of the patients. An aspect of particular interest, which indisputably has a crucial impact on the mechanics of the aorta, is its geometry. This is intuitively clear, and verified by multiple studies, such as stroke propensity analysis by Choi et al. [1] and the geometry-based analysis of new-onset heart failure by Beeche et al. [2]. Morphological studies have demonstrated that Type B aortic dissection patients exhibit significantly increased aortic arch dimensions, with longer arch lengths [3], whereas type A aortic dissection can potentially be predicted from abnormalities in the ascending aorta geometry [4]. Minderhoud et al.[5] argued that patients with a repaired aortic coarctation exhibit a smaller aortic arch, suggesting geometric features could be of importance in long-term risk assessment. As the outlined works verify, geometric parameters play a key role in understanding mechanics and patient health, though their effects remain underexplored.
Many works have studied the aortic geometry and its effects [6, 7, 8, 9, 10, 11]. However, a barrier that is ever-present lies in obtaining usable patient-specific models. Existing works usually make use of a small number of patient-specific geometries. This is completely understandable when one considers the difficulties and the user interaction required for obtaining computational models from patient-specific CT scans [12, 13, 14, 15]. Besides, obtaining the CT scans themselves already presents a challenge [16, 17].
@@
Imposing a parametrized model to generate (smooth) aortic geometries presents a challenging task. In our approach, we only require a single base model, containing a centerline which consists only of points and straight segments. Then, the aforementioned parameters are applied only to the centerline, which is an extremely simple object. While it is clear that a centerline may be extracted from the surface of a patient-specific aorta, going the other way around may not seem like a viable task. Here, we make use of convolution surfaces [46, 47], a graphics methodology to render surfaces based purely on centerline and radial data. Essentially, a small number of points, segments, and associated radii (order of magnitude of $20-30$) are enough to describe a smooth surface of an aorta. Thus, parametrizing the centerline and radius information enables the parametrization of the entire aorta. The base model and centerline are obtained from [14], where the centerline was computed based on the mesh contraction approach [48].
Even though the generated surfaces are useful, they are still insufficient to perform numerical simulations. The primary goal herein is to enable users to directly employ the finite element, finite volume, finite difference or any other mesh-based numerical method on high-quality meshes to effectively analyze the physical processes at work. With this objective in mind, it is necessary to produce volumetric discretizations of the generated aortic surfaces, i.e., meshes. The standard methodology for non-trivial domains such as aortas is unstructured meshing [49]. This is justified by the flexibility and robustness of the meshing procedure, as well as a multitude of off-the-shelf software tools [50, 51]. On the other hand, it is often worth investing extra effort into the generation of structured meshes [52, 14], as they offer certain advantages. Examples include exact control of the element count, straightforward formation of boundary layers, flow-oriented elements, and finally, the control of the local mesh structure. Hence, the resulting meshes have the same node placement independently of the geometry variation, greatly boosting their applicability for various applications such as machine learning or reduced order modeling. To realize the mesh generation, we make use of the block-structured approach presented in [14], extended in [53, 54]. It fits well with the approach described herein since the required input is exactly a convolution surface description of the domain.
Ultimately, this dataset presents the main step towards enabling geometry-based analyses of the aorta. The main barrier towards devising a machine learning or reduced order model that can reach conclusions on mechanical quantities purely from geometrical information, without performing new simulations, is in fact the lack of available simulation data. Performing numerical simulations on the high-quality meshes in this dataset provides exactly the key data necessary for such models.
@@ 2.1 Convolution surfaces
where $\Gamma:[0,l]\rightarrow\mathbb{R}^{3}$ denotes the parametrization of a centerline segment $[\mathbf{a},\,\mathbf{b}]$ of length $l$ defined as $\Gamma_{S}(s)=\mathbf{a}+\frac{s}{l}(\mathbf{b}-\mathbf{a}).$ The compactly supported kernel function $K:\mathbb{R}\rightarrow\mathbb{R}$ is defined as
@@
and the distance function $g:\mathbb{R}^{3}\times\mathbb{R}^{3}\rightarrow\mathbb{R}$ as $g(\mathbf{x},\mathbf{y})=\sqrt{\mathbf{x}^{T}\,G(\mathbf{y})\,\mathbf{x}}.$ The matrix $G$ is computed in an eigenvalue decomposition form $G=UDU^{T}.$ Following [47], $U$ is obtained by taking the rotation matrix that maps the vector $e_{1}=(1,0,0)$ to the tangent vector of the respective segment $\Gamma^{\prime}(s).$ The matrix $D$ captures the radius information using $D(s)=\textrm{diag}(\alpha(s),\beta(s),\gamma(s))$, controlling the linearity of the radius along the segment. The functions $\alpha,\beta\textrm{ and }\gamma$ are computed from the equation $\xi(s)=\big{(}\frac{l-s}{l}\xi_{0}^{-\frac{1}{2}}+\frac{s}{l}\xi_{1}{-\frac{1}{2}}\big{)}^{-2}$, where the distinct initial values $\xi_{0}$ and $\xi_{1}$ for each of the three functions are computed from the radiuses of the segment $r_{\mathbf{a}}$ and $r_{\mathbf{b}}$. This framework enables anisotropic convolution surfaces with non-circular cross-sections. We omit further details for brevity, as this construction mainly serves to prove the radius control as shown in [47].
@@ 2.2 Statistical geometric parametrization
Table 1: Aortic parameters and their probabilistic distributions. The parameters in group A are purposefully modified to create new aortic datasets. The parameters in group B are produced as a result of the newly generated aorta. The distribution parameters, e.g., $(a,b)$, are defined as mean and variance for a normal distribution, as location and scale for a Gumbel distribution, and as minimum and maximum values for a uniform distribution. Some of the parameter combinations produce invalid or unphysiological domains, causing the convolution surface to blend, in particular between the carotid arteries. In a minority of the cases, the meshing algorithm produces invalid elements. Thus, the final distributions are slightly affected. The units in the SynthAorta statistical distributions are preserved as in the literature.
Collected distribution SynthAorta distribution
Group Parameter Notation type parameters Unit Reference type parameters
A Ascending aorta radius $r_{\text{AA}}$ Normal $(14.08,2.59)$ mm [34, 35, 36, 37] Normal $(13.69,2.30)$
Aortic arch radius $r_{\text{Arch}}$ Normal $(13.85,3.16)$ mm [34, 36] Normal $(13.00,1.99)$
Descending aorta radius $r_{\text{DA}}$ Normal $(12.17,2.31)$ mm [38, 34, 36] Normal $(12.17,2.31)$
Curvature radius multiplier $\gamma_{\text{c}}$ Normal $(1.07,0.01)$- [7, 39, 11] Gumbel $(1.10,0.06)$
Descending aorta adjustment $\rho$- Uniform $(0,1)$
Nonplanar Fourier multiplier $\delta_{3}$- Normal $(1,0.09)$
Nonplanar Fourier multiplier $\delta_{4}$- Normal $(1,0.09)$
B Brachiocephalic artery (BCA) radius $r_{\mathrm{BCA}}$ Normal $(6.10,0.12)$ mm [40] Normal $(6.10,0.12)$
Common carotid artery (CCA) radius $r_{\mathrm{CCA}}$ Normal $(3.26,0.03)$ mm [42] Normal $(3.26,0.03)$
Left subclavian artery (LSA) radius $r_{\mathrm{LSA}}$ Normal $(5.42,0.07)$ mm [41] Normal $(5.42,0.07)$
Centerline curvature radius $R_{\text{c}}$ Normal $(39.30,3.92)$ mm [7, 39, 11] Gumbel $(40.41,2.40)$
@@ 2.2.1 Aortic radii
We use a Gaussian random field to create realistic radii samples of the three aortic segments, using a Karhunen-Loève (KL) expansion. It is used only for computational efficiency, since many realizations of the random field are needed. A typical SVD truncation happens in the background, yielding acceptable accuracy while greatly improving efficiency. This method helps us generate smooth, interconnected variations along the length of the aorta [55].
The correlation structure is manually adjusted to reflect the physiological connections between the different segments of the aorta, ensuring that the produced aortas maintain realistic anatomical characteristics. We determine the average and variability of the random field from weighted values obtained from medical literature. These variations introduce physiological differences within each realization. The mean and variance of the random field are defined as $\mu_{\text{w}}$ and $\sigma_{\text{w}}$ and are derived from weighted values reported in the medical literature, see Table 1.
@@ 2.2.3 Centerline curvature radius
For the rest of the section, we assume the base centerline has been mapped to a single best-fitting plane using a typical SVD-based procedure. This is done to keep the definition of the centerline curvature radius in line with existing work [11]. We define the point $\mathbf{A}$ in the ascending aorta, heuristically, as the point approximately closest to the bifurcation of the nearby pulmonary vessel, and consequentially $\mathbf{B}$ as the point on the centerline which is on the same axial plane as $\mathbf{A}$, the same as in [7] and [11], shown in Fig. 3a.
@@
Denoting the centerline points of the main aorta which lie between them: $\mathbf{A}=\mathbf{x_{1}},\mathbf{x_{2}},\dots,\mathbf{x_{n-1}},\mathbf{x_{n}}=\mathbf{B}$, the centerline curvature radius is determined as follows. For the initial patient-specific centerline, we determine the center point $\mathbf{x_{\text{c}}}=(x_{\text{c}},y_{\text{c}},z_{\text{c}}),$ and the radius $r$ of the circle by solving the following optimization problem
$\underset{\mathbf{x_{\text{c}}},r}{\min}\,\,{\sum_{i=1}^{n}\left(\left\|\mathbf{x_{i}}-\mathbf{x_{\text{c}}}\right\|-r\right)^{2}},$ (5)
@@ 2.2.4 Descending aorta adjustment
Let us consider the centerline points $\mathbf{x_{n}},\,\dots,\mathbf{x_{n+k}}$, starting from $\mathbf{x_{n}}=\mathbf{B}$ and ending at the main outlet point of the centerline $\mathbf{x_{n+k}}=\mathbf{x_{\text{outlet}}}$, with the connections between neighbouring points $\mathbf{x_{j}}$ and $\mathbf{x_{j+1}}$. Let $\Delta\mathbf{B}\in\mathbb{R}^{3}$ be the displacement of the point $\mathbf{B}$ resulting from a change in the centerline curvature radius $R_{\text{c}}$. We introduce a new parameter $\rho$ uniformly distributed between 00 and $1$, defining the displacement at the outlet point $\mathbf{x_{\text{outlet}}}$ as
@@
${\Delta\mathbf{x_{j}}=\Big{(}(1-\rho)\Big{(}1-\frac{\|\mathbf{B}-\mathbf{x_{j}}\|}{\|\mathbf{B}-\mathbf{x_{\text{outlet}}}\|}\Big{)}+\rho\Big{)}\Delta\mathbf{B},}$ (11)
where $j=n+1,\dots,n+k-1$. The displacement at $\mathbf{x_{j}}$ is a therefore a distance-based interpolation of displacements at $\mathbf{B}$ and $\mathbf{x_{\text{outlet}}}$. An example is illustrated in Fig. 3c.
@@ 2.2.5 Nonplanar displacement
$\mathbf{d}_{N}(\mathbf{x})=A_{0}+\sum_{n=1}^{N}\left(A_{n}\cos\left(nw{\|\mathbf{x}\|}\right)+B_{n}\sin\left(nw{\|\mathbf{x}\|}\right)\right),$ (12)
@@
$\hat{\mathbf{d}}_{2}^{(k)}(\mathbf{x})=A_{0}$ $+A_{1}\gamma_{1}\cos(w{\|\mathbf{x}\|})+B_{1}\gamma_{2}\sin(w{\|\mathbf{x}\|})$
$+A_{2}\delta_{3}\cos(2w{\|\mathbf{x}\|})+B_{2}\delta_{4}\sin(2w{\|\mathbf{x}\|}).$ (13)
@@ 2.3 Mesh generation
The algorithm builds a block-structure [57], a coarse domain subdivision, based on the input centerline. The block-structure generation is based on various centerline configurations occurring in the aorta. Then, a sub-mesh is generated in each of the blocks using transfinite interpolation [58]. In order to ensure conformity and matching between the blocks, points are first generated on the edges of the block-structure. Then, surface nodes are mapped to the domain’s surface using a gradient-based projection. Transfinite interpolation is then used to compute the face points using the already-computed edge points, subsequently applying the surface mapping again where necessary. Finally, the points in the block interiors are generated using the computed face and edge points via transfinite interpolation, completing the point generation procedure.
Once the submeshes inside the blocks are generated, the task of defining a global connectivity matrix simply means converting local node indices to global indices. Conformity is ensured during the procedure itself, making the connectivity matrix generation straightforward. More details on the procedure may be found in [14, 53].
@@
2.4 Vascular flow simulation setup
In order to demonstrate the application-readyness of the SynthAorta dataset, we employ the cases contained within the dataset for vascular flow simulation. As any other investigation starting from the dataset, we formulate a complete boundary value problem targeting the physics of interest. The setup considered herein is identical to [59], which considers the Navier–Stokes equations for incompressible flow of a generalized Newtonian fluid (see, e.g., [60, 61] for their numerical treatment),
$\frac{\partial}{\partial t}\mathbf{u}+\mathbf{u}\cdot\nabla\mathbf{u}-2\nabla\cdot\left(\nu\nabla^{\mathrm{S}}\mathbf{u}\right)+\nabla p$ $=\mathbf{0}$ in $\text{in }\Omega\times(0,T],$
$\nabla\cdot\mathbf{u}$ $=0$ in $\text{in }\Omega\times[0,T],$
with ${\nabla^{\text{S}}}(\cdot)\coloneqq\frac{1}{2}\left[\nabla(\cdot)+\nabla(\cdot)^{\top}\right]$, fluid velocity $\mathbf{u}$, pressure $p$ and variable kinematic viscosity $\nu$. The domain $\Omega$ is discretized for each individual case as contained within the SynthAorta dataset. The viscosity is governed by the Carreau model,
$\nu(\dot{\gamma})=\eta_{\infty}+(\eta_{0}-\eta_{\infty})\left[1+(\lambda\dot{\gamma})^{2}\right]^{\frac{n-1}{2}},\quad$
$\text{with}\quad\dot{\gamma}:=\sqrt{2\nabla^{\mathrm{S}}\mathbf{u}:\nabla^{\mathrm{S}}\mathbf{u}},$
capturing the shear-thinning behavior of blood with physiological fluid parameters taken from [24].
[…]
In the context of the present work, we show that the meshes contained in the SynthAorta dataset are of sufficient accuracy and quality for an application in vascular flow, to yield physiological results when combined with suitable simulation tools and data. To this end, we i) perform a convergence study on the relevant quantities of interest in vascular flows, and ii) provide the flow rates and pressures on each of the outlets together with the spatial average of the wall shear stress (WSS) in the descending aorta.
The simulations are performed using ExaDG [63] (see [64] and [65]). The software is based on deal.II [66] and its matrix-free infrastructure [67, 68], and supports $hp$-multigrid preconditioners based on [68, 69]. Within this setup, the construction of nested mesh hierarchies as provided within the SynthAorta dataset is vital to obtain a suitable geometric coarsening sequence within the $hp$-multigrid preconditioner, enabling sufficient resolution and large sampling sets.